Three point charges are arranged along x-axis as follows: charge q1 = 3.0 µC is
at the origin, and charge q2 = - 5.00 µC is at x = 0.200 m. Charge q3 = -8.00 µC.
Where is q3 located if the net force on q1 is 7.00 N in the -x-direction?
1
Expert's answer
2020-11-20T10:02:55-0500

A sketch of "q_3" between "q_1" and "q_2" pushes the Force"(F)" towards the left.

"F" from point "3" tends towards left.

"C=" "(-5\times -8)\over (0.200-x)^2"

"F" from point "1" also tends towards the left.

"\therefore C=" "(3.0\times 8)\over (x)^2"

"C=" "40\over (0.04-x^2)" "+" "24\over x^2" "=7.00"

"C=\mu^{3k}"